OpenStudy (anonymous):

I will give you a medal and be your fan To the nearest tenth of a percent, what is 5/7 written as a percent? A. 714.3% B. 71.4% C. 7.1% D. 0.7%

OpenStudy (whpalmer4):

You need to look at what you post when you copy/paste!

OpenStudy (anonymous):

@whpalmer4 I fixed it

OpenStudy (whpalmer4):

Can you tell me what 5/7 is as a decimal?

OpenStudy (anonymous):

0.71428571429

OpenStudy (whpalmer4):

Okay, if 1 = 100%, what do you think 0.714285714285... equals as a percentage?

OpenStudy (anonymous):

0.7?

OpenStudy (doc.brown):

\[\frac{0.714285714285}{1}\]

OpenStudy (whpalmer4):

no... \[\frac{0.714285}{1} = \frac{x\%}{100\%}\]

OpenStudy (whpalmer4):

0.714285 * 100% = 71.4285%, doesn't it?